OpenStudy (anonymous):

Fan and medal Solve using perfect square factoring patterns. x^2 – 10x + 25 = 9 A. {–4, 12} B. {4, –12} C. {–8, –2} D. {2, 8}

OpenStudy (anonymous):

Answer is D. You take 9 away from each side and then factor it our to be (x-8)(x-2). You then take the opposite of the numbers in the parenthesis, 2 and 8.

OpenStudy (cruffo):

or note that the left-hand side is a perfect square binomial, and factor. Then use the square root property to take the square root of both sides: \(x^2 - 10x +25 = 9\) \((x-5)^2 = 9\) \(x-5 = 3\) and \(x-5 = -3\) \(x = 8\) and \(x = 2\)